OpenStudy (anonymous):

What is the equation of the line that passes through the point (2, 5) and has a slope of 2? y = 2x + 1 y = 2x + 3 y = 2x - 8 y = 2x - 3

OpenStudy (shamil98):

(x,y) --> (2,5) m = slope = 2 y = mx + b input your values, solve for b(yintercept).

OpenStudy (anonymous):

so 5 = 2x + 5 ?

jimthompson5910 (jim_thompson5910):

m = 2 x = 2 y = 5 y = mx + b y = 2x + b ... plug in m = 2 (the given slope) 5 = 2*2 + b ... plug in (x,y) = (2,5) (the given point) now solve for b

OpenStudy (anonymous):

I'm still confused .. im not getting it

jimthompson5910 (jim_thompson5910):

do you see how I got 5 = 2*2 + b ?

OpenStudy (anonymous):

you plugged the numbers in ?

jimthompson5910 (jim_thompson5910):

yes, I plugged in the given slope ( m ) and the given point (x,y)

jimthompson5910 (jim_thompson5910):

you now need to solve for b in other words, you need to isolate b

OpenStudy (anonymous):

so how would i do that ?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

I'll get you started and let you finish 5 = 2*2 + b 5 = 4 + b 5 = b + 4 b + 4 = 5 b = ???

OpenStudy (anonymous):

1

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

thank you so much ! :)

jimthompson5910 (jim_thompson5910):

m = 2 (given slope) b = 1 (what you just solved for) y = mx + b y = 2x + 1

jimthompson5910 (jim_thompson5910):

sure thing